Transportation
media%2Fbb4%2Fbb4ca13e-9b0d-4189-a655-91Can correct the answer
4.A two-phase signal system is to be designed for an isolated intersection with a peak hour tactor 81 0.99. TIL lane volumes are: Phase A-550 veh/hr, Phase B 600 veh/hr, The total lost time L is 14 sec. Determine an appropriate cycle length for the intersection that will satisfy the following conditions: minimum cycle length-45 sec, maximum cycle length = 120 sec, maximum critical ve ratio (Xc)-0.85 and The saturation flow per lane for each phase is 1900 veh/hr.
Answer #1

Given:

XC = critical v/c ratio for the intersection = 0.85

S = Saturation flow rate =1900 veh/ hr

PHF = Peak Hour Flow = 0.95

L = total lost time per cycle = 14 sec

で) 7 c PHF v) = 49001 = 0.3047 550 for phasel, 600 for phase, 100.3324 0.95 Σ (-) = 0.3047 0.3324 = 0.6371 c C-L 0.850.6371 C

where,

C = cycle length (sec)

(v/c) = critical ratio = ratios of actual flow to saturation flow

Σ (v/c) = summation of the flow ratio